OpenStudy (anonymous):

Plz help, The length (l) of a rectangle is 8 mm longer than its width (w). Its perimeter (p) is more than 60 mm. Let w equal the width of the rectangle. P=2w+2l a. Write an expression for the length in terms of the width b. Use these expressions to write an inequality based on the given information. c. Solve the inequality, clearly indicating the width of the rectangle. Write your answer in a complete sentence.

OpenStudy (raden):

a). length (l) = w + 8 (known) b). perimeter > 60 2w + 2l > 60 because l = w + 8, so 2w + 2(w+8) > 60 2w + 2w + 16 > 60 4w + 16 > 60 c). 4w + 16 > 60 first, subtract by 16 on both sides, get 4w + 16 - 16 > 60 - 16 4w > 44 then divided by 4 on both sides, get 4w/4 > 44/4 w > 11 because w > 11 then length would be more than : length (l) = w + 8 = 19 length (l) > 19